A client has kefazolin 400 mg prescribed PO q 8h. Kefazolin 250 mg per tsp is available in oral suspension. How many mL...

  1. A client has kefazolin 400 mg prescribed PO q 8h. Kefazolin 250 mg per tsp is available in oral suspension. How many mL will the nurse administer per dose? How many mL would be administered if the concentration was kefazolin 250 mg per Tbsp?
Answer #1

Per dose amount to be administered = 400 mg

250 mg per tsp (5 ml)

400 mg = 5/250 * 400 = 8 ml

250 mg per Tbsp(15 ml)

400 mg = 15/250 * 400= 24 ml